# Symbolic Powers via Extension

### v2

@@ Abstract.
This article investigates under which conditions the symbolic powers of the extension of an ideal is the same as the extension of the symbolic powers. Our result generalizes the known scenarios. As an application, we prove formulas for the resurgence of sum of two homogeneous ideals in finitely generated ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is algebraically closed. Initially, these were known for ideals in polynomial rings.
@@ 1. Introduction
The symbolic powers are very well studied objects. The behavior of symbolic powers under homomorphism are also well studied. If $\phi:A\to B$ is a homomorphism of rings, then for any ideal $I$ in $A$ we use $IB$ to denote the extension of the ideal $I$ in $B$ and for any ideal $J$ in $B$ we use $J\cap A$ to denote the contraction of the ideal $J$ in $A$.
In [GHMN17, Proposition 3.8], it has been proved that if $\phi:A={\mathbbm{k}}[y_{1},\dots,y_{s}]\to B={\mathbbm{k}}[x_{0},\dots,x_{n}]$ is defined by $y_{i}\mapsto f_{i}$, with every subset of $c+1$ many $f_{i}$ forms $B$-regular sequence then $I^{(m)}B=\left(IB\right)^{(m)}$, where $I$ is the Stanley-Reisner ideal of a matroid on $[s]=\{1,\dots,s\}$ of dimension $s-1-c$. In [Ake17, Theorem 4], it has been proved that, if $\phi:A={\mathbbm{k}}[y_{0},\dots,y_{n}]\to B={\mathbbm{k}}[y_{0},\dots,y_{n}]$ is given by $y_{i}\to f_{i}$, where $f_{0},\dots,f_{n}$ is an $B$-regular sequence of homogeneous element of the same degree, then $(IB)^{(m)}=(I^{(m)})B$, where $I$ is the defining ideal of a 00-dimensional scheme. In [Wal18, Proposition 2.1], it has been proved that if $\phi:A\to B$ is faithfully flat and the associative primes of $I$ extends to associative primes of $IB$ then the symbolic powers of the extension is the extension of that of $I$.
Motivated by these, in this article, we investigate the following question.
@@
Let $A$ and $B$ be noetherian rings, and $I$ be an ideal of $A$. Let $\phi:A\to B$ be a ring homomorphism. When is $(IB)^{(m)}=I^{(m)}B$, where $m\geqslant 2$?
In [GHMN17, Proposition 3.8], [Ake17, Theorem 4], and [Wal18, Proposition 2.1], the key tool to prove symbolic power extends to symbolic power was that the morphisms were faithfully flat. We proved that one can relax faithfully flatness and prove something more general. We prove the following result extending the known situations where symbolic powers of the extension is the extension of the symbolic powers.
@@
In Remark 2.6, we discussed that our hypothesis is generalizes the known situations. Thus it covers more classes and will broader applications.
One of the central topic of interest in the study of symbolic powers is to find $i$ and $j$ for which $I^{(i)}\subset I^{j}$. In [ELS01, HH02, MS18], it has been proved that if $I$ is a homogeneous ideal of bigheight $h$ in a polynomial ring then $I^{(rh)}\subset I^{r}$ for all $r\geqslant 1$. In the other hand, one can investigate non-containment between the symbolic and ordinary powers of ideals. The resurgence, defined by Bocci and Harbourne [BH10], and the asymptotic resurgence, defined by Guardo, Harbourne, and Van Tuyl [GHVT13] measure such non-containment.
If $I$ is a homogeneous ideal in a finitely generated regular ${\mathbbm{k}}$-algebra domain $A$, where ${\mathbbm{k}}$ is perfect then the resurgence number, $\rho(I)$ and the asymptotic resurgence, $\rho_{a}(I)$ are defined as follows:
@@
If $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$ are ideals in polynomial rings ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ then the resurgence and the asymptotic resurgence numbers of $I+J\subset{\mathbbm{k}}[\underline{x},\underline{y}]$ attracted huge attention. Motivated by 1.1 we study the following question.
@@
Let $A$ and $B$ be regular noetherian rings, $I$ be an ideal of $A$, and $J$ be an ideal of $B$. What can be said about $\rho(I+J)$ and $\rho_{a}(I+J)$, where $I+J=IR+JR$, where $R=A\otimes B$?
The resurgence and the asymptotic resurgence of the sum of ideals $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$, where ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ are polynomial rings, are studied in [BHJT21, KNT24]. As an application of Theorem 2.5 and Proposition 3.3, we extend those results ideals in standard graded finitely algebra domains over algebraically closed fields.
@@
Let $I\subset A$ and $J\subset B$ be homogeneous ideals in finitely generated regular ${\mathbbm{k}}$-algebra domains $A$ and $B$, where ${\mathbbm{k}}$ is algebraically closed. Set $I+J=IR+JR$, where $R=A\otimes B$. Then the following hold .
@@
    (Theorem 4.5) $\rho_{a}(I+J)=\overline{\rho}_{a}(I+J)=\overline{\rho}(I+J)={\max}\{\rho_{a}(I),\rho_{a}(J)\}=\max\{\rho_{a}(IR),\rho_{a}(JR)\}.$
@@
    (Theorem 4.6) $\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$ Furthermore, if $\max\{\rho(I),\,\rho(J)\}\geqslant 2\min\{\rho(I),\,\rho(J)\}$, then $\rho(I+J)=\max\{\rho(I),\,\rho(J)\}$.
The paper is outlined as follows. In Section 2, we prove our main results regarding the relation symbolic powers of an ideal with its extension. In Section 3 we study prime extension property and its applications. In Section 4, we apply results from Section 2 and Section 3 we extend results regarding the sum of ideals in polynomial rings to finitely generated algebra domains over algebraically closed fields.
@@ Acknowledgments.
The authors are thankful to Tài Huy Hà for suggesting the problem and valuable comments. This is a part of the second author’s Ph.D. thesis. We are also thankful to Souvik Dey and Hop D. Nguyen for some helpful discussions.
2. Symbolic Powers via flat extension
In this section we prove that with specific conditions, the symbolic powers of an extension is the extension of the symbolic powers. Before proceeding to the main theorem we need to prove the following lemma regarding flat morphisms.
@@ 2. Symbolic Powers via flat extension
Proof.
The first part can be found in a standard book, e.g., Dummit and Foote. But we provide the proof for completion. After tensoring the exact sequence $0\rightarrow I\rightarrow R\rightarrow R/I\rightarrow 0$ by $M$, we get $I\otimes M\rightarrow R\otimes M\rightarrow(R/I)\otimes M\rightarrow 0$. It’s a well-known fact that $R\otimes_{R}M\cong M$, then it induces a module homomorphism $\phi$ from $I\otimes M$ to $M$ defined by $\phi:I\otimes M\to M,\phi(r\otimes m)=rm$. Now the image of $\phi$ is $IM$. Since $R\otimes_{R}M\cong M$, the kernel of induced homomorphism from $M$ to $(R/I)\otimes M$ is exactly the image of $\phi$, then by isomorphism theorem, we have $M/IM\cong(R/I)\otimes M$. Since $M$ is a flat, $R$-module we get the following exact sequence: $0\rightarrow I\otimes M\rightarrow R\otimes M\rightarrow(R/I)\otimes M\rightarrow 0$. Following the previous result, we immediately have $I\otimes_{R}M\cong IM$. ∎
Now we are ready to prove one of the key lemmas.
@@
Now we prove that $\psi_{2}$ is also well-defined. Indeed, $c\notin\mathfrak{p}$ and $d\notin\mathfrak{p}B$, then $\phi(c)d\notin\mathfrak{p}B$ because $\phi(A)\cap\mathfrak{p}B=\phi(\mathfrak{p})$. To see this, let $\mathfrak{p}=\mathfrak{q}\cap A$ for an ideal $\mathfrak{q}$ in $B$. From this we observe that $\mathfrak{p}B\cap A=\left(\left(\mathfrak{q}\cap A\right)B\right)\cap A=\mathfrak{q}\cap A={\mathfrak{p}}$. This implies that $\phi(A)\cap\mathfrak{p}B=\phi(\mathfrak{p})$. Note that $\phi(A)\cap\mathfrak{p}B\supset\phi(\mathfrak{p})$. Let $y\in\phi(A)\cap{\mathfrak{p}}B$. Then $y=\phi(x)$ for some $x\in A$. We show that $x\in{\mathfrak{p}}$. Now since $y\in{\mathfrak{p}}B$ then $x\in{\mathfrak{p}}B\cap A={\mathfrak{p}}$.
Notice that $\psi_{1}$ and $\psi_{2}$ are inverses to each other.
Now we show that $\psi_{1}$ is an $A$-algebra homomorphism, then we are done. Verifying the map commutes with the multiplication is straightforward, we only verify the addition here. Let $\frac{a_{1}\otimes b_{1}}{c_{1}\otimes d_{1}}$ and $\frac{a_{2}\otimes b_{2}}{c_{2}\otimes d_{2}}$ be elements in $((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)$. Now
@@
[Mat89, Theorem 23.2] Let $\phi:A\to B$ be a homomorphism of Noetherian rings, and Let $E$ be an $A$-module and $G$ a $B$-module. If $G$ is flat over $A$, then we have the following
@@
Note that by [HJKN23, Lemma 2.2] if ${\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})$ then ${\mathfrak{p}}\subset{\mathfrak{q}}$ for some ${\mathfrak{q}}\in\operatorname{Ass}(I)$. Thus we get $\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})}{\mathfrak{p}}\subset\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}$ for all $n\geqslant 1$. Since $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$ then using Theorem 2.3 we get
$\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(I^{(n)}B)}{\mathfrak{q}}=\bigcup_{\begin{subarray}{c}{\mathfrak{q}}\in\operatorname{Ass}({\mathfrak{p}}B),\\ {\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)\end{subarray}}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})}{\mathfrak{p}}B\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B,\text{ as }\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)}\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}.$
The following proves that if a flat morphism extends associative primes to associative primes and each associative prime is a contraction of some ideal then symbolic powers of the extension of an ideal is the extension of the symbolic powers of the ideal.
@@
First we check that as in Lemma 2.4, $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$. Using the definition of symbolic powers of an ideal and assumption, we get
@@
Now apply Lemma 2.2, we have $B_{\mathfrak{p}B}\cong A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$. Then by Lemma 2.1, we have the following:
@@
$=(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$
(*) $\cong(I^{n}A_{\mathfrak{p}}\cap A)B_{\mathfrak{p}B}\cap B$
Now we prove the last isomorphism. Observe that $B=A\otimes_{A}B\subset A\otimes_{A}B_{\mathfrak{p}B}$. Applying a property from [Mat89, section 7], which states that if $M_{2}$ is flat over $M_{1}$ and $M_{3}$ is flat over $M_{2}$, then $M_{3}$ is flat over $M_{1}$. Here we consider $M_{3}$ as $B_{\mathfrak{p}B}$, $M_{2}$ as $B$, and $M_{1}$ as $A$, then we can see $B_{\mathfrak{p}B}$ is flat over $A$. Now by [Mat89, Theorem 7.4(I)], we consider $I^{n}A_{\mathfrak{p}}$ and $A$ as submodules of $A_{\mathfrak{p}}$ as $A$-modules. Then $\left(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}\right)\cap\left(A\otimes_{A}B_{\mathfrak{p}B}\right)=(I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B}\cong(I^{n}A_{\mathfrak{p}}\cap A)B_{\mathfrak{p}B}$, then we intersect it with $B$ as $A\otimes_{A}B\cong B$. Recall that $I^{(n)}=\bigcap_{\mathfrak{p}\in\operatorname{Ass}_{A}(I)}(I^{n}A_{\mathfrak{p}}\cap A)$. Therefore we have:
$(IB)^{(n)}=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}(IB)^{n}B_{{\mathfrak{p}}B}\cap B$
[…]
@@
First we prove equality 2.1. Since $\phi$ is flat, then using Lemma 2.1, we get
$I^{(n)}B_{{\mathfrak{p}}B}=\left(\cap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}(I^{n}A_{\mathfrak{p}}^{\prime}\cap A)\right)B_{{\mathfrak{p}}B}=\left(\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}(I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)\right)\otimes_{A}B_{{\mathfrak{p}}B}$
$=\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}((I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)\otimes_{A}B_{{\mathfrak{p}}B})=\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}((I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)B_{{\mathfrak{p}}B})=(I^{n}A_{\mathfrak{p}}\cap A)B_{{\mathfrak{p}}B}.$
@@
The proof of equality 2.3 is similar to [Gri21, Remark 2.8]. We add the proof for completion. Note that $I^{(n)}B_{W}\cap B=\{r\in B~{}|~{}rs\in I^{(n)}\text{ for some }s\in W\}$ and thus $I^{(n)}B_{W}\cap B\subset\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\left(\left(I^{(n)}B_{{\mathfrak{p}}B}\right)\cap B\right)$. Now let $r\in\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\left(\left(I^{(n)}B_{{\mathfrak{p}}B}\right)\cap B\right)$. Then for each ${\mathfrak{p}}B$ there is $s\not\in{\mathfrak{p}}B$ such that $rs\in I^{(n)}B$. Thus $I^{(n)}B:r$ is not contained in any ${\mathfrak{p}}B\in\operatorname{Ass}(IB)$. Since the associated primes are finite then by Prime avoidance theorem
@@
Thus there is a $s\in W$ such that $sr\in I^{(n)}B$. Hence we have the equality as sets.
Again the equality 2.4 is straight forward. Lastly, we prove equality 2.5. Now from [AM94, Proposition 3.11(ii)]

### v3

@@ Abstract.
This article investigates under which conditions symbolic powers of the extension of an ideal are the same as the extension of the symbolic powers. As an application, we prove formulas for the resurgence of sum of two homogeneous ideals in finitely generated graded $\mathbbm{k}$}-algebras which are domains, where $\mathbbm{k}$} is algebraically closed. Initially, these were known for homogeneous ideals in polynomial rings.
@@ 1. Introduction
The symbolic powers are very well-studied objects for commutative algebra and algebraic geometry, especially for the multiplicity theory. We are interested in the behavior of symbolic powers under ring homomorphisms. If $\phi:A\to B$ is a homomorphism of rings, then for an ideal $I$ in $A$, we use $IB$ to denote the extension of the ideal $I$ in $B$, and for an ideal $J$ in $B$, we use $J\cap A$ to denote the contraction of the ideal $J$ in $A$. In this article, we primarily focus on the following question.
@@
Let $A$ and $B$ be Noetherian rings, and $I$ be an ideal of $A$. Let $\phi:A\to B$ be a ring homomorphism. When is $(IB)^{(n)}=I^{(n)}B$, where $m\geqslant 1$?
If the map $\phi$ is faithfully flat, then the Question 1.1 has been well studied, for examples check, [GHMNmatroid, Theorem 3.6], [Akessesh, Theorem 4], and [walker2018uniform, Proposition 2.1]. Our result generalizes from faithfully flat map to flat map, which is the following:
@@
We also give examples that satisfy assumptions in Theorem 2.5 but not under the faithfully flat map, such as Example 2.8 and Example 2.9. Moreover, In Example 2.7, we show that the condition that associated primes are contractions is necessary.
One of the central topics of symbolic powers is the containment between the symbolic and ordinary powers of ideals. The containment problem asks to find $m$ and $r$ for which $I^{(m)}\subset I^{r}$. In [EinLazarsfeldSmith, HochsterHuneke, MaScwede], it has been proved that if $I$ is a proper ideal of big height $h$ (the largest height of associated primes of $I$) in a regular ring containing a field, then $I^{(rh)}\subset I^{r}$ for all $r\geqslant 1$. After this celebrated result, one can keep investigating exact numbers for containments. The resurgence, defined by Bocci and Harbourne [bocci2010comparing], and the asymptotic resurgence, defined by Guardo, Harbourne, and Van Tuyl [guardo2013asymptotic] measure containment in more refined settings.
If $I$ is a proper ideal in a Notherian ring $R$, $\rho(I)$ and the asymptotic resurgence, $\rho_{a}(I)$ are defined as follows:
@@
Motivated by Question 1.1 and resurgence numbers, we study the following question.
@@
Let $A$ and $B$ be Notherian and regular $\mathbbm{k}$}-algebras, II be an ideal of $A$, and $J$ be an ideal of $B$. What can be said about $\rho(I+J)$ and $\rho_{a}(I+J)$, where $I+J=IR+JR$ and $R=A\otimes_{{\mathbbm{k}}}B$?
The resurgence and the asymptotic resurgence of the sum of ideals $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$, where ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ are polynomial rings, are studied in [fiberprojective, sharpbound]. As an application of Theorem 2.5, we extend those results ideals in standard graded finitely generated $\mathbbm{k}$}-algebras which are domains, and $\mathbbm{k}$} is algebraically closed.
@@
Let $I\subset A$ and $J\subset B$ be homogeneous ideals in finitely generated graded regular $\mathbbm{k}$}-algebra domains $A$ and $B$, where $\mathbbm{k}$} is algebraically closed. Set $I+J=IR+JR$, where $R=A\otimes_{{\mathbbm{k}}}B$. Then the following hold.
@@
    (Theorem 3.5) $\rho_{a}(I+J)={\max}\{\rho_{a}(I),\rho_{a}(J)\}=\max\{\rho_{a}(IR),\rho_{a}(JR)\}$, and also
    $\rho_{a}(I+J)=\overline{\rho}_{a}(I+J)=\overline{\rho}(I+J)$, where $\bar{\rho}$ and $\bar{\rho_{\alpha}}$ are resurgence (asymptotic resurgence respectively) of integral closure.
@@
    (Theorem 3.6) $\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$
    If $\max\{\rho(I),\,\rho(J)\}\geqslant 2\min\{\rho(I),\,\rho(J)\}$, then $\rho(I+J)=\max\{\rho(I),\,\rho(J)\}$.
The paper is outlined as follows. In Section 2, we prove our main result regarding the relation symbolic powers of an ideal with its extension, and we also study prime extension property and its applications. In Section 3, we apply results from Section 2 and we extend results regarding the sum of ideals in polynomial rings to finitely generated algebra domains over algebraically closed fields.
@@ Acknowledgments.
The authors are thankful to Tài Huy Hà for suggesting the problem and valuable comments. We are also thankful to Souvik Dey and Hop D. Nguyen for some helpful discussions.
2. Symbolic Powers and Resurgence via flat extension
In this section we will show our main theorem (Theorem 2.5). Before proceeding to the main theorem we provide some useful lemmas. The first lemma is very well-known and we state it without proof.
@@ 2. Symbolic Powers and Resurgence via flat extension
The next result is also a standard one. Since we could not find a resource, we provide a complete proof.
@@
We prove that $\psi_{2}$ is also well-defined. Indeed, $c\notin\mathfrak{p}$ and $d\notin\mathfrak{p}B$, then $\phi(c)d\notin\mathfrak{p}B$ because $\phi(A)\cap\mathfrak{p}B=\phi(\mathfrak{p})$. To see this, let $\mathfrak{p}=\mathfrak{q}\cap A$ for an ideal $\mathfrak{q}$ in $B$. From this we observe that $\mathfrak{p}B\cap A=\left(\left(\mathfrak{q}\cap A\right)B\right)\cap A=\mathfrak{q}\cap A={\mathfrak{p}}$. This implies that $\phi(A)\cap\mathfrak{p}B=\phi(\mathfrak{p})$. Note that $\phi(A)\cap\mathfrak{p}B\supset\phi(\mathfrak{p})$. Let $y\in\phi(A)\cap{\mathfrak{p}}B$. Then $y=\phi(x)$ for some $x\in A$. We show that $x\in{\mathfrak{p}}$. Now since $y\in{\mathfrak{p}}B$ then $x\in{\mathfrak{p}}B\cap A={\mathfrak{p}}$.
Notice that $\psi_{1}$ and $\psi_{2}$ are inverses to each other. Thereofre we just need to show that $\psi_{1}$ is an $A$-algebra homomorphism, then we are done. Verifying the map commutes with the multiplication is straightforward, we only verify the addition here. Let $\frac{a_{1}\otimes b_{1}}{c_{1}\otimes d_{1}}$ and $\frac{a_{2}\otimes b_{2}}{c_{2}\otimes d_{2}}$ be elements in $((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)$. Then
@@
[matsumura1989commutative, Theorem 23.2] Let $\phi:A\to B$ be a homomorphism of Noetherian rings, and let $E$ be an $A$-module and $G$ a $B$-module. If $G$ is flat over $A$, then we have the following
@@
Note that by [ha2023binomial, Lemma 2.2] if ${\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})$ then ${\mathfrak{p}}\in\operatorname{Ass}(I^{n})$ and ${\mathfrak{p}}\subset{\mathfrak{q}}$ for some ${\mathfrak{q}}\in\operatorname{Ass}(I)$. Thus we get $\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})}{\mathfrak{p}}\subset\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}$ for all $n\geqslant 1$. Since $\operatorname{Ass}^{*}(IB)=\{{\mathfrak{p}}B~|~{\mathfrak{p}}\in\operatorname{Ass}^{*}(I)\}$ then using Theorem 2.3 we get
$\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(I^{(n)}B)}{\mathfrak{q}}=\bigcup_{\begin{subarray}{c}{\mathfrak{q}}\in\operatorname{Ass}({\mathfrak{p}}B),\\ {\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)\end{subarray}}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})}{\mathfrak{p}}B\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B,\text{ as }\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)}{\mathfrak{p}}\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}.$
We are ready to prove the main theorem of this section. The theorem states that if a flat morphism extends associative primes to associative primes and each associative prime is a contraction of some ideal, then symbolic powers of the extension of an ideal is the extension of the symbolic powers of the ideal.
@@
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}^{*}(IB)=\{{\mathfrak{p}}B~|~\text{ for all }{\mathfrak{p}}\in\operatorname{Ass}^{*}(I)\}$, and every associated prime of $I$ is a contraction of some ideal in $B$. Then $(IB)^{(n)}=I^{(n)}B$, for all $n\geqslant 1$. Furthermore, if $I^{l}$ is a contraction of some ideals for all $\ell\geqslant 1$, then $I^{(m)}\subset I^{r}$, if and only if $IB^{(m)}\subset IB^{r}$, for all $m,r\geqslant 1$, then we get $\rho(IB)=\rho(I)$ and $\rho_{a}(IB)=\rho(I)$.
@@
First we check that as in Lemma 2.4, $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~|~{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$. Using the definition of symbolic powers of an ideal and the assumption, we get
@@
Applying Lemma 2.2, we have $B_{\mathfrak{p}B}\cong A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$. Then by Lemma 2.1, we have the following:
@@
Next we want to show that $((I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)=(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$.
It is known that $A\otimes_{A}B\subset A\otimes_{A}B_{\mathfrak{p}B}$ if $S=B\backslash\mathfrak{p}B$ has no zero divisors. Suppose $S$ has no zero divisors, then applying a property from [matsumura1989commutative, section 7], which states that if $M_{2}$ is flat over $M_{1}$ and $M_{3}$ is flat over $M_{2}$, then $M_{3}$ is flat over $M_{1}$. Here we consider $M_{3}$ as $B_{\mathfrak{p}B}$, $M_{2}$ as $B$, and $M_{1}$ as $A$, then we can see $B_{\mathfrak{p}B}$ is flat over $A$. By [matsumura1989commutative, Theorem 7.4(I)], we consider $I^{n}A_{\mathfrak{p}}$ and $A$ as submodules of $A_{\mathfrak{p}}$ as $A$-modules, then $\left(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}\right)\cap\left(A\otimes_{A}B_{\mathfrak{p}B}\right)=(I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B}$. Therefore, intersecting it with $A\otimes_{A}B$ will not change the original intersection.
If $S$ has zero divisors, then $B\backslash B_{\mathfrak{p}B}$ has exactly the same set of zero divisors in $S$, denoted as $S^{\prime}$. Note that any $a\otimes b\in I^{n}A_{\mathfrak{p}}\otimes_{A}S^{\prime}$ is always 0 in $I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$, and similarly for $a\otimes b\in A\otimes_{A}S^{\prime}$ in $A\otimes_{A}B_{\mathfrak{p}B}$. Observe that these 0’s in $A\otimes_{A}B_{\mathfrak{p}B}$ are 0’s in $(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$, so we can always intersect $I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$ with $A\otimes_{A}B_{\mathfrak{p}B}$ first, then $A\otimes_{A}B$, which leads to the same equality. Therefore $(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)=((I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$ as we desire.
Applying lemma 2.1 again on $((I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$, we will have our second isomorphism $((I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)\cong(I^{n}A_{p}\cap A)B_{{\mathfrak{p}}B}\cap B$. Note that lemma 2.1 is used previously for the “backward” direction for the first isomorphism, and now applying it again for the “forward” direction for our second isomorphism, so we will have a new equality (two isomorphisms give an identity map), which is $(IB)^{n}B_{\mathfrak{p}B}\cap B=(I^{n}A_{p}\cap A)B_{{\mathfrak{p}}B}\cap B$. Then the following equality is given:
$(IB)^{(n)}=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}(IB)^{n}B_{{\mathfrak{p}}B}\cap B=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}((I^{n}A_{p}\cap A)B_{{\mathfrak{p}}B})\cap B$
We want to prove the following labeled equalities:
$(IB)^{(n)}=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}((I^{n}A_{p}\cap A)B_{{\mathfrak{p}}B})\cap B$
@@
First we prove equality 2.1. Since $\phi$ is flat, and extension commutes with intersection under flat morphism, we get
$I^{(n)}B_{{\mathfrak{p}}B}=\left(\cap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}(I^{n}A_{\mathfrak{p}}^{\prime}\cap A)\right)B_{{\mathfrak{p}}B}=\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}((I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)B_{{\mathfrak{p}}B})=(I^{n}A_{\mathfrak{p}}\cap A)B_{{\mathfrak{p}}B}.$
@@
The proof of equality 2.3 is similar to [grifo2021symbolic, Remark 2.8]. We add the proof for completion. Note that $I^{(n)}B_{W}\cap B=\{r\in B~|~rs\in I^{(n)}\text{ for some }s\in W\}$ and thus
$\displaystyle I^{(n)}B_{W}\cap B\subset\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\left(\left(I^{(n)}B_{{\mathfrak{p}}B}\right)\cap B\right).$
Now let $r\in\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\left(\left(I^{(n)}B_{{\mathfrak{p}}B}\right)\cap B\right)$. Then for each ${\mathfrak{p}}B$ there is $s\not\in{\mathfrak{p}}B$ such that $rs\in I^{(n)}B$. Thus $I^{(n)}B:r$ is not contained in any ${\mathfrak{p}}B\in\operatorname{Ass}(IB)$. Since the associated primes are finite then by Prime avoidance theorem
@@
Then there is a $s\in W$ such that $sr\in I^{(n)}B$. Hence we have the equality as sets.
Again the equality 2.4 is straight forward. Lastly, we prove equality 2.5. Now from [atiyah1994introduction, Proposition 3.11(ii)]